OpenStudy (anonymous):

What is (sqrt(2sinx) + sqrt(2sinx)) simplified to?

OpenStudy (anonymous):



OpenStudy (anonymous):

Oops, I meant 2sin(x)

OpenStudy (unklerhaukus):

\[\sqrt{2\sin x} +\sqrt{2\sin x}=2\sqrt{2\sin x}\]

OpenStudy (unklerhaukus):

i dont know if you can simplify more than that

sam (.sam.):

Middle steps \[\sqrt{2\sin x} +\sqrt{2\sin x} \\ \\ \sqrt{2}\sqrt{sinx}+\sqrt{2}\sqrt{sinx} \\ \\ \sqrt{2}(2\sqrt{sinx}) \\ \\ 2\sqrt{2}\sqrt{sinx} \\ \\ 2\sqrt{2sinx}\]